OpenStudy (anonymous):

Medal and fan Reduce the fraction to lowest terms. Do not use spaces in your answer. And please explain!!

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{ 5r -5s }{ 5r +5s } = \frac{ 5*(r-s) }{ 5*(r+s) } = \frac{ (r-s) }{ (r+s) }\]